# Leptogenesis and Neutrino Masses Via Pseudo-Dirac Gauginos

### v1

@@ III.1 Bino oscillations
and $\Gamma(\psi_{\tilde{B}}\to h\ell)$ is the partial decay width for an initial $|\psi_{\tilde{B}}\rangle$ state to decay into a lepton state. Setting $\Delta\Gamma/(2\Gamma_{\tilde{B}})=0$, in the limit $r\ll 1$ and $m\gg r\Gamma_{\tilde{B}}$, we have the following approximation,
where $x\equiv\Delta m/\Gamma_{\tilde{B}}\simeq 2m/\Gamma_{\tilde{B}}$. Note that $CP$ violation is maximized for $x\sim 1$. Since this approximation is valid in the parameter regime we consider, we will use eq. 22 in the calculation of the baryon asymmetry later.
III.2 Other bino interactions
Unlike generic RH neutrinos in various seesaw models, the bino in this model has copious interactions with the SM particles. Elastic scatterings of the bino with the SM fermions in the plasma and bino–antibino annihilations affect the bino–antibino oscillations in the early universe, and hence the baryon asymmetry that will be generated through their decays. Assuming a common sfermion mass, $M_{\rm sf}$, the relevant effective interaction Lagrangian, after integrating out heavy sfermions, can be written as Hsieh (2008),
$-\mathcal{L}_{\rm int}$ $=\frac{2g_{Y}^{2}}{M_{\rm sf}^{2}}\left(Y_{L}^{2}\bar{\psi}_{\tilde{B}}^{C}P_{L}F\,\bar{F}P_{R}\psi_{\tilde{B}}^{C}+Y_{R}^{2}\bar{\psi}_{\tilde{B}}P_{R}F\,\bar{F}P_{L}\psi_{\tilde{B}}\right),$ (23)
where $F^{T}=(f_{L},~{}\bar{f}^{\dagger}_{R})$ and $Y_{L,R}$ are the hypercharges for the left- and right-handed SM fermions, $f_{L}$ and $f_{R}$, respectively. We can re-write the above Lagrangian, using Fierz identities, to read,
@@ III.2 Other bino interactions
$\begin{split}\mathcal{L}_{\rm int}&=\mathcal{L}_{\rm blind}+\mathcal{L}_{\rm sens}\,,\\ \mathcal{L}_{\rm blind}&=\frac{g_{Y}^{2}}{4M_{\rm sf}^{2}}\bar{\psi}_{\tilde{B}}\gamma_{\mu}\gamma_{5}\psi_{\tilde{B}}\bar{F}\gamma^{\mu}(g_{V}-g_{A}\gamma_{5})F\,,\\ \mathcal{L}_{\rm sens}&=\frac{g_{Y}^{2}}{4M_{\rm sf}^{2}}\bar{\psi}_{\tilde{B}}\gamma_{\mu}\psi_{\tilde{B}}\bar{F}\gamma^{\mu}(g_{V}-g_{A}\gamma_{5})F\,.\\ \end{split}$ (25)
$\mathcal{L}_{\rm blind}$ is symmetric under $\psi_{\tilde{B}}\to\psi_{\tilde{B}}^{C}$ and $\mathcal{L}_{\rm sens}$ is antisymmetric. (Note that $F$ is kept unchanged.) The thermally-averaged annihilation and scattering cross-sections for flavor-blind and flavor-sensitive interactions are,
$\begin{split}\langle\sigma\,v\rangle_{\rm sens}^{\rm ann}&=\frac{C\,M_{\tilde{B}}^{2}}{4M_{\rm sf}^{4}}\,,\quad\langle\sigma\,v\rangle_{\rm blind}^{\rm ann}=\frac{C\,M_{\tilde{B}}^{2}}{4M_{\rm sf}^{4}}\frac{1}{z}\,,\\ \langle\sigma\,v\rangle_{\rm sens}^{\rm scat}&=\frac{C\,M_{\tilde{B}}^{2}}{4M_{\rm sf}^{4}}\frac{3z^{2}+6z+7}{6(z^{4}+z^{3})}\simeq\frac{C\,M_{\tilde{B}}^{2}}{4M_{\rm sf}^{4}}\frac{1}{2z^{2}}\,,\\ \langle\sigma\,v\rangle_{\rm blind}^{\rm scat}&=\frac{C\,M_{\tilde{B}}^{2}}{4M_{\rm sf}^{4}}\frac{9z^{2}+6z+2}{6(z^{4}+z^{3})}\simeq\frac{C\,M_{\tilde{B}}^{2}}{4M_{\rm sf}^{4}}\frac{3}{2z^{2}}\,,\end{split}$ (26)
@@
$C=\frac{g_{Y}^{4}}{4\pi}\left(\sum_{i={\rm leptons}}\left(Y_{Ri}^{4}+Y_{Li}^{4}\right)+3\sum_{i={\rm quarks}}\left(Y_{Ri}^{4}+Y_{Li}^{4}\right)\right)\,.$
The sums are over all the quark and lepton flavors. We take all SM fermions, including the top quark, to be massless since we will consider $M_{\tilde{B}}\sim\mathcal{O}({\rm TeV})$.
@@ III.3 Oscillations in the early universe
The $CP$ violation necessary for generating the BAU requires that bino–antibino oscillations take place in the early universe. We follow the density matrix formalism described in Tulin et al. (2012) in order to incorporate the oscillation dynamics in the early universe. This method was developed for neutrino oscillations in medium and the early universe Dolgov (1981); Sigl and Raffelt (1993) and was later applied to asymmetric dark matter models Buckley and Profumo (2012); Cirelli et al. (2012); Tulin et al. (2012).
@@
$\begin{split}zH\frac{d\mathbb{Y}}{dz}=-i(\mathbb{HY}-\mathbb{YH^{\dagger}})-\sum_{+,-}\frac{\Gamma_{\pm}}{2}\left[O_{\pm}\left[O_{\pm},\mathbb{Y}\right]\right]\\ -\sum_{+,-}s\langle\sigma v\rangle_{\pm}\left(\frac{1}{2}\{\mathbb{Y},O_{\pm}\mathbb{\bar{Y}}O_{\pm}\}-Y_{\rm eq}^{2}\right),\end{split}$ (27)
where $z=M_{\tilde{B}}/T,s=2\pi^{2}g_{*}T^{3}/45$ is the entropy density, $\mathbb{H}$ is the Hamiltonian given in eq. 13 and
@@
is the equilibrium abundance for $z>1$. $O_{\pm}={\rm diag}(1,\pm 1)$ and the subscript $\pm$ differentiate between interaction terms that are symmetric (+)(+)( + ) or antisymmetic ()(-)( - ) under the transformation $\psi_{\tilde{B}}\to\psi_{\tilde{B}}^{C}$. For binos, both types of interactions are present. (See eq. 25.) Note that the above equation is only valid for non-relativistic particles, ignoring effects $\mathcal{O}(p^{2}/M^{2})$, where $p\sim T$. (The Hamiltonian is given for $\mathbb{p}=0$.) Hence, we only use this equation for $z=M_{\tilde{B}}/T>1$. The density matrices are,
$\mathbb{Y}=\begin{pmatrix}Y_{\psi_{\tilde{B}}}&Y_{\psi_{\tilde{B}}\psi_{\tilde{B}}^{C}}\\ Y_{\psi_{\tilde{B}}^{C}\psi_{\tilde{B}}}&Y_{\psi_{\tilde{B}}^{C}}\end{pmatrix}\,,~{}~{}\mathbb{\bar{Y}}=\begin{pmatrix}Y_{\psi_{\tilde{B}}^{C}}&Y_{\psi_{\tilde{B}}^{C}\psi_{\tilde{B}}}\\ Y_{\psi_{\tilde{B}}\psi_{\tilde{B}}^{C}}&Y_{\psi_{\tilde{B}}}\end{pmatrix}\,,$
@@
In eq. 27 the first term describes oscillations in vacuum in the presence of $CP$ violation and decays. The second term accounts for scatterings, where $\Gamma_{\pm}$ is the scattering rate. This term vanishes for scatterings that are symmetric under $\psi_{\tilde{B}}\to\psi_{\tilde{B}}^{C}$. The last term is for annihilations and $\langle\sigma v\rangle_{\pm}$ is the thermally-averaged annihilation cross section.
@@
$zH\frac{d\Sigma}{dz}=$ $-\Gamma_{\tilde{B}}\Sigma+2\Gamma_{\tilde{B}}r(i\sin\phi\,\Xi-\cos\phi\,\Pi)$
@@
$zH\frac{d\Delta}{dz}=$ $-\Gamma_{\tilde{B}}\,\Delta+2im\,\Xi\,,$ (30)
@@
where the relevant cross sections are given in eq. 26. As a representative example, in fig. 2 we plot the bino and antibino yields for $M_{\tilde{B}}=1$ TeV, $M_{\rm sf}=50$ TeV, $\Lambda_{M}=4\times 10^{7}$ TeV and $m=10^{-4}$ eV. It can be seen in fig. 2 that the oscillations start around $z\sim 12$, corresponding to $T\sim 80$ GeV, which is below $T_{\rm sph}\simeq 130$ GeV. This set of parameters was chosen to showcase the oscillations with $CP$ violation. In the next section, we will focus on larger $m$ values so that bino oscillations start before sphalerons shut off.
Figure 2: The bino (orange) and antibino (purple) abundances for an initial asymmetry of zero. For $CP$ violating parameter we take $r=0.2$ and $\sin\phi=1$. With $M_{\tilde{B}}=1$ TeV and $\Lambda_{M}=4\times 10^{7}$ TeV, we have $\Gamma_{\tilde{B}}\simeq 2.5\times 10^{-5}$ eV.
Let us estimate when the oscillations commence. Coherent bino–antibino oscillations can only start after the Hubble rate falls below the oscillation frequency, $\omega_{\rm osc}=\Delta m>H(T)$. However, bino interactions with the SM plasma tend to hinder oscillations as well. This phenomenon is an example of the quantum Zeno effect Misra and Sudarshan (1977), in which scatterings and annihilations can act as measurements on a quantum system. This effect can be analytically understood by looking at the evolution equations for $\Delta(z)$ and $\Xi(z)$ above. For an approximate analysis, we can omit $\Gamma_{\tilde{B}}$ as decays and oscillations are independent. We can then solve for $\Xi$ from the second equation in section III.3 and use this in the third equation. After a change of variables with $y=z^{2}$, we get
@@
where $\tilde{\Gamma}(y)=2\,\Gamma_{\rm sens}^{\rm scat}+s\,\Sigma\,\langle\sigma v\rangle_{\rm blind}^{\rm ann}$. (Note that the combination $yH$ is a constant.) This equation describes a damped harmonic oscillator where the damping term drops over time. Oscillations occur only in an under-damped system. In the existence of interactions, the start of oscillations is when the following under-damping requirement is satisfied,
@@
$\frac{m^{2}}{y^{2}H^{2}}$ $>\left(\frac{\tilde{\Gamma}(y)}{4yH}\right)^{2}\Rightarrow 4m>\left(2\,\Gamma_{\rm sens}^{\rm scat}+s\,\Sigma\,\langle\sigma v\rangle_{\rm blind}^{\rm ann}\right)_{z=z_{\rm osc}}.$ (33)
@@
If the Majorana mass is large enough, oscillations could start while the bino is relativistic. However, our treatment in eq. 27 does not capture this regime. To ensure the validity of our analysis, we limit our focus to the parameter space where oscillations begin at temperatures below the bino mass $T_{\rm osc}<M_{\tilde{B}}$. Furthermore, it is preferred for the bino to start oscillating before it decays to maximize the CP violation. Since we focus on bino decays that occur before the sphalerons shut off, $T>T_{\rm sph}=130~{}$GeV D’Onofrio et al. (2014), we also require $z_{\rm osc}<z_{\rm sph}=M_{\tilde{B}}/T_{\rm sph}$. These requirements translate into a condition on the Majorana mass $m$ range,
$\tilde{\Gamma}(z=1)>4m>\tilde{\Gamma}(z=z_{\rm sph})\,.$ (34)
@@
This region is shown in fig. 3, assuming a thermal bino distribution for the annihilation rate [^3] For $z=1$, the bino is close to equilibrium and for $z=z_{\rm sph}$, scattering rate largely dominates over the annihilation rate.. For the Dirac bino masses we consider, $\mathcal{O}({\rm TeV})$, the Majorana masses need to be larger than $\sim 10^{-3}$ eV so that the oscillations are not delayed below the sphaleron temperature. The upper bound, beyond which our non-relativistic analysis does not apply, depends on the bino mass more strongly. The interactions are suppressed with larger $M_{\rm sf}$. Hence, oscillations can start earlier, limiting the region of validity. We choose $M_{\rm sf}=50$ TeV as a benchmark value. These conditions could be relaxed with a much more detailed analysis. However, as we show below, a large parameter space where the baryon asymmetry can be generated already exists even with these limitations.
@@
Figure 3: The parameter space, $M_{\tilde{B}}$ vs the Majorana mass $m$, for which bino oscillations start between $T=M_{\tilde{B}}$ and $T_{\rm sph}=130$ GeV. (See eq. 34 and the accompanying text.) Solid/green, dashed/blue and dashed-dotted/purple lines are for $M_{\rm sf}=25,50,{\rm~{}and~{}}100$ TeV, respectively. In the bottom-shaded region bino oscillations start after sphalerons turn off whereas in the upper-shaded regions, oscillations are expected to start while bino is relativistic, $T>M_{\tilde{B}}$.
@@ IV The Baryon Asymmetry Of the Universe
As the binos oscillate and decay in the early universe, the resulting lepton asymmetry can be quantified using the $CP$-violating parameter in eq. 21:
@@
where $\Delta_{L}=Y_{L}-Y_{\bar{L}}$. Here we include a washout term from the inverse decays which is especially important for $M_{\rm sf}\lesssim 25$ TeV as the bino interactions are large enough to keep them close to equilibrium before they decay. We ignore another washout term coming from $\Delta_{L}=2$ scatterings, $h\ell\to h\bar{\ell}$ through bino exchange, since these are highly suppressed by a factor of $|M_{\tilde{B}}\tilde{m}/\Lambda_{M}^{2}|^{2}$.
The lepton asymmetry is partially converted to a baryon asymmetry through the electroweak sphalerons in the SM. Sphaleron processes conserve $B-L$ while violating $B+L$. By doing a detailed balance of chemical potentials in various electroweak interactions, one finds,
@@
where $n_{H}$ is the number of Higgs doublets. (See, e.g., Davidson et al. (2008) for a review.) We use a minimally supersymmetric model where $n_{H}=2$. This conversion ceases when the sphalerons fall out of equilibrium at a temperature of $T_{\rm sph}=130~{}$GeV D’Onofrio et al. (2014). Thus, the resulting baryon asymmetry can be determined as,
@@
Figure 4: Lines of constant $\Delta_{B}$ for $\Lambda_{M}=2\times 10^{7}$ TeV, $M_{\rm sf}=50$ TeV, $r=0.2$ and $\phi=\pi/2$. (The observed value is $\sim 8\times 10^{-11}$.) The largest asymmetry produced is slightly smaller than $10^{-6}$ (not shown), between the two red lines in the bottom-right corner corresponding to $\Delta_{B}=10^{-7}$.
@@
Figure 5: The baryon asymmetry as a function of the Majorana mass $m$ for $M_{\tilde{B}}=1$ TeV. The horizontal dashed line represents the observed yield of BAU, $\Delta B\sim 8.7\times 10^{-11}$. We take $r\sin\phi=0.2$ and $\Lambda_{M}=2\times 10^{7}$ TeV. Solid, dashed, and dashed-dotted lines are for $M_{\rm sf}=25,~{}50,{~{}\rm and~{}}100~{}\rm TeV$, respectively. The lines are cut off using the condition in eq. 34.
@@
In fig. 4 we show the lines of constant $\Delta_{B}$ for varying $M_{\tilde{B}}$ and $m$, setting $\Lambda_{M}=2\times 10^{7}$ TeV, $M_{\rm sf}=50$ TeV, $r=0.2$ and $\sin\phi=1$. The baryon asymmetry can exceed the observed value in a significant portion of the parameter space, reaching up to $10^{-7}$, shown as the red line. For a given value of $M_{\tilde{B}}$, the baryon asymmetry first increases with decreasing $m$ as $x=2m/\Gamma_{\tilde{B}}$ approaches 1. When $x$ drops below 1, the asymmetry gets smaller again. This is why there are two lines corresponding to $\Delta_{B}\simeq 10^{-7}$. In between these lines, the largest value of the asymmetry is slightly below $10^{-6}$. As can be seen in fig. 5, the baryon asymmetry can slightly exceed $10^{-6}$ with $M_{\rm sf}=100$ TeV. This allows for smaller values of $r$ and $\phi$, or more specifically, smaller values of the combination $r\,\sin\phi$, which is the relevant combination for $\varepsilon$ (as seen in eq. 21). For constant Majorana mass $m$, heavier binos produce a larger asymmetry because the lifetime $\tau_{B}\sim\Lambda_{M}^{2}/M_{\tilde{B}}^{3}$ is shorter and the bino decays while its number density is larger. Smaller sfermion masses result in a smaller asymmetry due to larger annihilation/pair-production cross-sections, which keeps the bino in equilibrium longer. This feature can be seen in fig. 5 where we plot the baryon asymmetry for $M_{\tilde{B}}=1$ TeV and $M_{\rm sf}=25,50,100$ TeV. These interactions cease to be relevant for $M_{\rm sf}\gtrsim 500~{}$ TeV.

### v2

@@ III.2 Other bino interactions
where $F^{T}=(f_{L},~\bar{f}^{\dagger}_{R})$ and $Y_{L,R}$ are the hypercharges for the left- and right-handed SM fermions, $f_{L}$ and $f_{R}$, respectively. We can rewrite the above Lagrangian, using Fierz identities, to read,
@@
[…]
$\mathcal{L}_{\rm blind}$ is symmetric under $\psi_{\tilde{B}}\to\psi_{\tilde{B}}^{C}$ and $\mathcal{L}_{\rm sens}$ is antisymmetric. (Note that $F$ is kept unchanged.) The thermally-averaged annihilation and scattering cross-sections for flavor-blind and flavor-sensitive interactions are
[…]
@@
$C=\frac{g_{Y}^{4}}{4\pi}\left(\sum_{i={\rm leptons}}\left(Y_{Ri}^{4}+Y_{Li}^{4}\right)+3\hskip-5.69046pt\sum_{i={\rm quarks}}\left(Y_{Ri}^{4}+Y_{Li}^{4}\right)\right).$
The sums run over all the quark and lepton flavors. We take all SM fermions, including the top quark, to be massless, since we consider $M_{\tilde{B}}\sim\mathcal{O}({\rm TeV})$.
@@ III.3 Oscillations in the early universe
The $CP$ violation necessary for generating the BAU requires that bino–antibino oscillations take place in the early universe. We follow the density matrix formalism described in Tulin:2012re in order to incorporate the oscillation dynamics in the early universe. This method was developed for neutrino oscillations in the medium and the early universe Dolgov:1980cq; Sigl:1993ctk and was later applied to asymmetric dark matter models Buckley:2011ye; Cirelli:2011ac; Tulin:2012re.
@@
$\begin{split}zH\frac{d\textbf{Y}}{dz}=-i(\textbf{HY}-{\textbf{Y}\textbf{H}^{\dagger}})-\sum_{+,-}\frac{\Gamma_{\pm}}{2}\left[O_{\pm}\left[O_{\pm},\textbf{Y}\right]\right]\\ -\sum_{+,-}s\langle\sigma v\rangle_{\pm}\left(\frac{1}{2}\{\textbf{Y},O_{\pm}{\bar{\textbf{Y}}}O_{\pm}\}-Y_{\rm eq}^{2}\right),\end{split}$ (27)
where $z=M_{\tilde{B}}/T,s=2\pi^{2}g_{*}T^{3}/45$ is the entropy density, H is the Hamiltonian given in eq. 13 and
@@
is the equilibrium abundance for $z>1$. $O_{\pm}={\rm diag}(1,\pm 1)$ and the subscript $\pm$ differentiate between interaction terms that are symmetric $(+)$ or antisymmetic ()(-) under the transformation $\psi_{\tilde{B}}\to\psi_{\tilde{B}}^{C}$. For binos, both types of interactions are present. (See eq. 25.) Note that the above equation is only valid for non-relativistic particles, ignoring effects $\mathcal{O}(p^{2}/M^{2})$, where $p\sim T$. (The Hamiltonian is given for $\textbf{p}=0$.) Hence, we only use this equation for $z=M_{\tilde{B}}/T>1$. The density matrices are
$\textbf{Y}=\begin{pmatrix}Y_{\psi_{\tilde{B}}}&Y_{\psi_{\tilde{B}}\psi_{\tilde{B}}^{C}}\\ Y_{\psi_{\tilde{B}}^{C}\psi_{\tilde{B}}}&Y_{\psi_{\tilde{B}}^{C}}\end{pmatrix}\,,~~{\bar{\textbf{Y}}}=\begin{pmatrix}Y_{\psi_{\tilde{B}}^{C}}&Y_{\psi_{\tilde{B}}^{C}\psi_{\tilde{B}}}\\ Y_{\psi_{\tilde{B}}\psi_{\tilde{B}}^{C}}&Y_{\psi_{\tilde{B}}}\end{pmatrix}\,$,
@@
In eq. 27, the first term describes oscillations in vacuum in the presence of $CP$ violation and decays. The second term accounts for scatterings, where $\Gamma_{\pm}$ is the scattering rate. This term vanishes for scatterings that are symmetric under $\psi_{\tilde{B}}\to\psi_{\tilde{B}}^{C}$. The last term is for annihilations, and $\langle\sigma v\rangle_{\pm}$ is the thermally-averaged annihilation cross section.
@@
$zH\frac{d\Sigma}{dz}=$ $-\Gamma_{\tilde{B}}\Sigma+2\Gamma_{\tilde{B}}r(i\sin\phi\,\Xi-\cos\phi\,\Pi)$ (30)
@@
$zH\frac{d\Delta}{dz}=$ $-\Gamma_{\tilde{B}}\,\Delta+2im\,\Xi\,$,
@@
where the relevant cross sections are given in eq. 26. As a representative example, in fig. 2 we plot the bino and antibino yields for $M_{\tilde{B}}=1$ TeV, $M_{\rm sf}=50$ TeV, $\Lambda_{M}=4\times 10^{7}$ TeV and $m=10^{-4}$ eV. It can be seen in fig. 2 that the oscillations start around $z\sim 12$, corresponding to $T\sim 80$ GeV, which is below $T_{\rm sph}\simeq 130$ GeV. This set of parameters was chosen to showcase the oscillations with $CP$ violation. In the next section, we will focus on larger $m$ values so that bino oscillations start before the sphalerons shut off.
Figure 2: The bino (orange) and antibino (purple) abundances for an initial asymmetry of zero. For CPCP violating parameter we take $r=0.2$ and $\sin\phi=1$. With $M_{\tilde{B}}=1$ TeV and $\Lambda_{M}=4\times 10^{7}$ TeV, we have $\Gamma_{\tilde{B}}\simeq 2.5\times 10^{-5}$ eV.
Let us estimate when the oscillations commence. Coherent bino–antibino oscillations can only start after the Hubble rate falls below the oscillation frequency, $\omega_{\rm osc}=\Delta m>H(T)$. However, bino interactions with the SM plasma tend to hinder oscillations as well. This phenomenon is an example of the quantum Zeno effect 10.1063/1.523304, in which scatterings and annihilations can act as measurements on a quantum system. This effect can be analytically understood by looking at the evolution equations for $\Delta(z)$ and $\Xi(z)$ above. For an approximate analysis, we can omit $\Gamma_{\tilde{B}}$ as decays and oscillations are independent. We can then solve for $\Xi$ from the second equation in eq. 30 and use this in the third equation. After a change of variables with $y=z^{2}$, we get
@@
where $\tilde{\Gamma}(y)=2\,\Gamma_{\rm sens}^{\rm scat}+s\,\Sigma\,\langle\sigma v\rangle_{\rm blind}^{\rm ann}$. (Note that the combination $yH$ is a constant.) This equation describes a damped harmonic oscillator in which the damping term drops over time. Oscillations occur only in an under-damped system. In the existence of interactions, the start of oscillations is when the following under-damping requirement is satisfied,
@@
$\frac{m^{2}}{y^{2}H^{2}}$ $>\hskip-1.42271pt\left(\hskip-0.56917pt\frac{\tilde{\Gamma}(y)}{4yH}\hskip-0.56917pt\right)^{2}\hskip-3.41418pt\Rightarrow 4m>\hskip-1.42271pt\left(2\,\Gamma_{\rm sens}^{\rm scat}+s\,\Sigma\,\langle\sigma v\rangle_{\rm blind}^{\rm ann}\right)_{z=z_{\rm osc}}\hskip-1.42271pt.$ (33)
@@
If the Majorana mass is large enough, oscillations could start while the bino is relativistic. However, our treatment in eq. 27 does not capture this regime. To ensure the validity of our analysis, we limit our focus to the parameter space where oscillations begin at temperatures below the bino mass $T_{\rm osc}<M_{\tilde{B}}$. Furthermore, it is preferable for the bino to start oscillating before it decays to maximize the CP violation. (Since we focus on bino decays that occur before the sphalerons shut off, $T>T_{\rm sph}=130$~GeV PhysRevLett.113.141602, we automatically satisfy $z_{\rm osc}<z_{\rm sph}=M_{\tilde{B}}/T_{\rm sph}$.) These requirements translate into a condition on the Majorana mass $m$ range,
$\tilde{\Gamma}(z=1)>4m>\tilde{\Gamma}(z=z_{\rm dec})\,$, (34)
@@
$z_{\rm dec}$ is when the bino decays. This region is shown in fig. 3, assuming a thermal bino distribution for the annihilation rate [^3] For $z=1$, the bino is close to equilibrium and for $z=z_{\rm dec}\gg 1$, scattering rate largely dominates over the annihilation rate. . In the next section, we will focus on two benchmark scenarios, depicted in the upper and lower plots of fig. 3. In the first one (upper), we fix $\Lambda_{M}=4\times 10^{7}$ GeV, in which larger bino masses will result in shorter lifetimes. In the second scenario (lower), we fix the bino width such that $T_{\rm dec}\simeq 160$ GeV, which requires larger messenger scales for heavier binos, up to $O(10^{8}~{\rm TeV})$. For the Dirac bino masses we consider $\mathcal{O}({\rm TeV})$, the Majorana masses need to be larger than $\sim 10^{-3}$ eV so that the oscillations are not delayed after the bino starts decaying. In the upper shaded regions, bino is expected to start oscillating while in equilibrium, and our non-relativistic analysis does not apply. (We note that in order to produce the correct baryon asymmetry, we will be far from this near-relativistic region.) The interactions are suppressed with larger $M_{\rm sf}$. Hence, oscillations can start earlier, limiting the region of validity. We will choose $M_{\rm sf}=100$ TeV as a benchmark value in the next section. These conditions could be relaxed with a much more detailed analysis. However, as we show below, a large parameter space where the baryon asymmetry can be generated already exists, even with these limitations.
@@
Figure 3: The parameter space, $M_{\tilde{B}}$ vs the Majorana mass $m$, for which bino oscillations start between $T=M_{\tilde{B}}$ and $T_{\rm dec}$ for $\Lambda_{M}=4\times 10^{7}$ TeV with varying $T_{\rm dec}$ (upper) and $T_{\rm dec}\simeq 162.5$ GeV with varying $\Lambda_{M}$ (lower). (See eq. 34 and the accompanying text.) Dashed/blue and dashed-dotted/purple lines are for $M_{\rm sf}=50,{\rm~and~}100$ TeV, respectively. In the bottom-shaded region, bino oscillations start after sphalerons turn off, whereas in the upper-shaded regions, oscillations are expected to start while bino is relativistic, $T>M_{\tilde{B}}$.
@@ IV The Baryon Asymmetry Of the Universe
As the binos oscillate and decay in the early universe, the resulting lepton asymmetry can be quantified using the $CP$-violating parameter in eq. 21444The asymmetry parameter $\epsilon$ arises from the interference between decays occurring with and without bino–antibino oscillations, as in the neutral $B$-meson system.:
@@
where $\Delta_{L}=Y_{L}-Y_{\bar{L}}$. (The asymmetry parameter $\epsilon$ is time-integrated, which is a good approximation for the regime we consider where $x\simeq 1-1000$.) Here, we include a washout term from the inverse decays, which is especially important for $M_{\rm sf}\lesssim 25$ TeV, as the bino interactions are large enough to keep them close to equilibrium before they decay. We ignore another washout term coming from $\Delta_{L}=2$ scatterings, $h\ell\to h\bar{\ell}$ through bino exchange, since these are highly suppressed by a factor of $|M_{\tilde{B}}\tilde{m}/\Lambda_{M}^{2}|^{2}$.
The lepton asymmetry is partially converted to a baryon asymmetry through the electroweak sphalerons in the SM. Sphaleron processes conserve BLB-L while violating $B+L$. By doing a detailed balance of chemical potentials in various electroweak interactions, one finds
@@
where $n_{H}$ is the number of Higgs doublets. (See, e.g., Davidson_2008 for a review.) We use a minimally supersymmetric model where $n_{H}=2$. This conversion ceases when the sphalerons fall out of equilibrium at a temperature of $T_{\rm sph}=130$~GeV PhysRevLett.113.141602. Thus, the resulting baryon asymmetry can be determined as
@@
Figure 4: Lines of constant $\Delta_{B}$ for $\Lambda_{M}=4\times 10^{7}$ TeV (upper) and $T_{\rm dec}\simeq 162$ GeV (lower) with $M_{\rm sf}=100$ TeV, $r=0.2$ and $\phi=\pi/2$. The observed value is $\sim 8\times 10^{-11}$.
@@
Figure 5: The baryon asymmetry as a function of the Majorana mass $m$ for benchmark parameters $M_{\tilde{B}}=2$ TeV and $r\sin\phi=0.2$. The horizontal dashed line represents the observed yield of BAU, $\Delta B\sim 8.7\times 10^{-11}$. Solid/purple lines are for fixed $\Lambda_{M}=4\times 10^{7}$ TeV, and dotdashed/blue lines are for fixed bino lifetime $T_{\rm dec}=1.25\,T_{\rm sph}\simeq 160{\rm\,GeV}$, corresponding to $\Lambda_{M}\simeq 10^{8}$ TeV for $M_{\rm sf}=50,~75,{~\rm and~}100~\rm TeV$. The lines are cut off using the condition in eq. 34.
@@
In fig. 4 we show the lines of constant $\Delta_{B}$ for varying $M_{\tilde{B}}$ and $m$, setting $M_{\rm sf}=100$ TeV, $r=0.2$ and $\sin\phi=1$ for two scenarios. In the first case, shown in the upper figure, we set $\Lambda_{M}=4\times 10^{7}$ TeV and in the second scenario, shown in the lower figure, we fix the bino decays to happen at $T_{\rm dec}\simeq 160$ GeV. In both cases, the baryon asymmetry can exceed the observed value in a significant portion of the parameter space, reaching up to $10^{-7}$, shown as the red lines. For a given value of $M_{\tilde{B}}$, the baryon asymmetry increases with decreasing $m$ as $x=2m/\Gamma_{\tilde{B}}$ approaches to 1. This allows for smaller values of $r$ and $\phi$, or more specifically, smaller values of the combination $r\,\sin\phi$, which is the relevant combination for $\varepsilon$ (as seen in eq. 21). For a constant Majorana mass $m$ and constant $\Lambda_{M}$, heavier binos produce a larger asymmetry because the lifetime $\tau_{B}\sim\Lambda_{M}^{2}/M_{\tilde{B}}^{3}$ is shorter and the bino decays while its number density is larger. Smaller sfermion masses result in a smaller asymmetry due to larger annihilation/pair-production cross-sections, which keeps the bino in equilibrium longer. This feature can be seen in fig. 5 where we plot the baryon asymmetry for $M_{\tilde{B}}=2$ TeV and $M_{\rm sf}=50,75,100$ TeV. These interactions cease to be relevant for $M_{\rm sf}\gtrsim 500$~ TeV.
At this point, let us situate our work in the realm of resonant leptogenesis Pilaftsis:2003gt; Pilaftsis:2004xx; Pilaftsis:2005rv; Chauhan:2021xus; Klaric:2021cpi; King:2024idj; Das:2024gua and ARS leptogenesis Akhmedov:1998qx; Drewes:2017zyw; Drewes:2021nqr; Caputo:2018zky; Baumholzer:2018sfb. Arguably, the most notable difference of our model is the production of heavy neutral leptons, which are binos in our case. Binos can always be produced copiously in the early universe through their non-neutrino interactions with the SM. After their production, their out-of-equilibrium decay is governed by their mixing with neutrinos, as in other leptogenesis scenarios. Hence, unlike the freeze-in scenario of ARS leptogenesis, we are always in the freeze-out regime as in resonant leptogenesis. An additional effect that arises from considering binos as right-handed neutrinos is their scattering off the SM plasma through sfermion exchange. These interactions delay bino–antibino oscillations, which we use to limit our parameter space to the regime where binos oscillate only when non-relativistic. Lastly, in the hybrid inverse seesaw mechanism, the wino interactions are the main source of neutrino masses, while the long-lived bino is responsible for creating the matter–antimatter asymmetry. As discussed in Section II, this sequestering is assumed to ensure that the bino decays at $T<M_{\tilde{B}}$ instead of the strong washout regime.